Calculation of equilibrium constant from pk,s: coloure - Colores ОН + NaHCO3 = + H.CO use benzoic acids pka Key
Answer #1

Ooh + NaHCO: T O LONE + H2CO3 pka = 6.35 pka = 4.14 Keq = 10 opka =10(6.35.4.19) = 102-16 Opka = difference in pka s = 144.5